OpenStudy (anonymous):

write an equation of the line that passes through the points (-1,4) and (-5,2)

OpenStudy (muhammad_nauman_umair):

suppose first point is (x1,y1) and second point is(x2,y2) then formula (y-y1)/(y2-y1)=(x-x1)/(x2-x1)

OpenStudy (muhammad_nauman_umair):

now u can solve

OpenStudy (anonymous):

y= x/2 + 9/2

OpenStudy (anonymous):

slope = (-1) x coefficient/y coefficient now put in formula: (y-y1)=slope(x-x1)